OpenStudy (anonymous):

the function f(x)=8x/8x+5 is one to one. find its inverse f^-1(x)

OpenStudy (anonymous):

yikes you gotta solve \[x=\frac{8y}{8y+5}\] for \(y\) any ideas ?

OpenStudy (anonymous):

i can walk you through it if you like

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok but maybe first we solve \[\frac{8y}{8y+5}=7\] so we can understand what is going on

OpenStudy (anonymous):

\[8y=7(8y+5)\\ 8y=56y+35\\ -48y=35\\ y=-\frac{35}{48}\]

OpenStudy (anonymous):

use the same steps, only without numbers, just variables \(x\) and \(y\)

OpenStudy (anonymous):

\[x=\frac{8y}{8y+5}\] multiply both sides by \(8y+5\) get \[x(8y+5)=8y\] multiply out on the left get \[8xy+5x=8y\] add \(5x\) to both sides and subtract \(8y\) get \[8xy-8y=-5x\] factor out the \(y\) \[(8x-8)y=-5x\] divide \[y=\frac{-5x}{8x-8}\]

OpenStudy (anonymous):

i said something wrong, should be "subtract \(5x\) from both sides" but the answer is right i think

OpenStudy (anonymous):

wanna check it?

OpenStudy (mosaic):

Another way to do this is to take the reciprocal as the first step: 1/x = 1 + 5/8y

OpenStudy (anonymous):

yeah i guess what would work too in this case, since the numerator has only one term

OpenStudy (anonymous):

@babijenny you ok with this?